OpenStudy (anonymous):

Let f be the function given by f(x) = e^x/2 Find the first three nonzero term and the general term around 0 for (e^x/2)-1)/x Use it to find g'(2) and use it to show that

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} n/ 4(n+1)! = 1/4\]

OpenStudy (anonymous):

I actually found the first three nonzero terms but don't understand the second part.

OpenStudy (anonymous):

Help please!

OpenStudy (anonymous):

Is it \[ \frac{ e^x} 2 \]

OpenStudy (anonymous):

What is g(x)?

OpenStudy (anonymous):

g(x) is the ((e^x/2 )-1)/x

OpenStudy (anonymous):

take the derivative of \(\frac{e^{\frac{x}{2}}-1}{x}\) i think is the first step,then replace x by 2

OpenStudy (anonymous):

derivative evaluated at 2 will give you \(\frac{1}{4}\) which means that the expansion evaluated at two (that infinite sum you have in the question) must also be \(\frac{1}{4}\)

OpenStudy (anonymous):

How do we do that?

OpenStudy (anonymous):

FOr the last part.

OpenStudy (anonymous):

how do you do what?

OpenStudy (anonymous):

Make the expansion evaluated at 2 = 1/4. If I plug it in, I get like some crazy fraction. How do you make the sum = 1/4